A car , starting from rest , accelerates in a straight line path at a constant rate of 2.5 m/s^2. How far will the car travel in 12 seconds? A.180 m B.120 m C.30 m D.15 m E.4.8 m


1
Expert's answer
2021-06-11T16:59:07-0400
"d=v_0t+\dfrac{1}{2}at^2,""d=0+\dfrac{1}{2}\cdot2.5\ \dfrac{m}{s^2}\cdot(12\ s)^2=180\ m."

Therefore, the correct answer is (A)180 m.